OpenStudy (anonymous):

Cot^2(x)-1/(1+cot^2(x))=1-2sin^2(x) Verify Identity

OpenStudy (anonymous):

I cannot verify this ID. Stuck somewhere in the middle.

OpenStudy (anonymous):

I end up with Cos^2(x)/Sin^2(x) - Sin^2(x) = Cos^2(x) + Sin^2(X) (or 1)

OpenStudy (raden):

use the identity Cot^2(x) = csc^2(x) - 1

OpenStudy (raden):

Cot^2(x)-1/(1+cot^2(x)) = (csc^2(x) - 1-1)/(1 + csc^2(x) - 1) = (csc^2(x) - 2)/( csc^2(x) ) = csc^2(x)/csc^2(x) - 2/csc^2(x) = 1 - 2sin^2(x) does that make sense ?

OpenStudy (anonymous):

Let me take a look at what I have and I'll get right back to you.

OpenStudy (anonymous):

yes it does. Thank you. I kept trying to change everything from Cot's to Cos's and Sin's right off the bat.

OpenStudy (raden):

you're welcome ;)